how many three digit numbers have only even digits?
How many ways are there to rearrange the letters in the word aptitude,if the first and last letter must each be a vowel?

Found 2 solutions by galactus, Edwin McCravy:
Answer by galactus(183)

There are 4 even digits: 2,4,6,8
There are 4 choices for the first place, the second place, and the third place
4*4*4=64 three digit numbers that have all even digits.

~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
There are 4!/2!=12 ways to place the vowels on the ends:
ae,ai,au,ea,ei,eu,ia,ie,iu,ua,ue,ui
Place them and arrange the 6 inside letters in 6!/2! ways.
We divide by 2 because of the repeated 't'.
So, the number of ways is 12*(6!/2!)=4320 ways

Answer by Edwin McCravy(20054)


We can choose the 1st digit as 2,4,6, or 8.  That's 4 ways.
We can choose the 2nd digit as 0,2,4,6, or 8.  That's 5 ways.
We can choose the 3rd digit as 0,2,4,6, or 8.  That's 5 ways.

Answer 4*5*5 = 100 three-digit numbers with only even digits.

Edwin